OpenStudy (anonymous):

hey im stuggling on a question, can someone please help how do you get from x(t) = e^(-t/3)u(t) to X(s) = 3/(3s + 1) x

OpenStudy (anonymous):

laplace transform?

OpenStudy (anonymous):

yeah but can not seem to get the answer. Help? I am really struggling x

OpenStudy (anonymous):

\[\large{L(e^{-{\frac{t}{3}}}u(t))= \frac{1}{s+\frac{1}{3}}}= \frac{3}{3s+1}\]

OpenStudy (amistre64):

hmm, not to adept with these, but it looks like maybe: \[-\frac{1}{3}t-st=-t(s+\frac{1}{3})\] how you get to Lanas I aint got no clue :) but afterwards you can simplify it by mutliplying at all by 3/3

OpenStudy (anonymous):

but u cant have t's in X(s)?

OpenStudy (amistre64):

right, which is the part im not to adept at ;)

OpenStudy (anonymous):

How would you show this from first principles? x

OpenStudy (amistre64):

ack!!

OpenStudy (amistre64):

pauls online notes has some good material

OpenStudy (anonymous):

im not sure i did it right either, what i did was laplace(e^(-t/3)) ...

OpenStudy (anonymous):

\[e^{-\alpha t}u(t)= \frac{1}{s+ \alpha}\]